OpenStudy (anonymous):

3√2/3 - 2√150 + 3√3/2

OpenStudy (anonymous):

Is it this? \[3\sqrt{\frac{2}{3}} - 2\sqrt{150} + 3\sqrt{\frac{3}{2}}\]

OpenStudy (anonymous):

haha yes!! thanks :)

OpenStudy (anonymous):

Ok sweet. Step 1: rationalize the denominators. Step 2: Find the least common denominator. Step 3: combine terms as much as you can

OpenStudy (anonymous):

Do you know how to rationalize the denominators?

OpenStudy (anonymous):

not really